OpenStudy (anonymous):

Solve the given problem. Melanie is 2 years older than twice her brother Bill’s age. If Bill is 5 years old, how old is Melanie? A. 4 B. 10 C. 12 D. 14

OpenStudy (anonymous):

12

OpenStudy (anonymous):

let age of melaaince=x then 2*5+2=x =>x=10+2=12. x=12

OpenStudy (camerondoherty):

Since bill is 5 you multilply 5 times 2 since it is double the age and add two \[5 \times 2 + 2=?\]

OpenStudy (anonymous):

@keepsecure93